OpenStudy (tanabugg):

Write an equation in​ point-slope form of the line that passes through the given​ points, then write the equation in​ slope-intercept form. ​(−9​,7​), ​(9,1) What is the​ point-slope form of the equation of the​ line?

563blackghost (563blackghost):

First find the slope... \(\huge\bf{\frac{y_{2}-y_{1}}{x_{2}-x_{1}} \rightarrow \frac{11-7}{9+9}=slope}\)

OpenStudy (tanabugg):

Well I already did that eith the rise and run but idk how to make an equation

OpenStudy (tanabugg):

@563blackghost

OpenStudy (tanabugg):

6/18 = 1/3 Can you look over the question again

563blackghost (563blackghost):

We are already given the first points which is `(-9,7)` we would input it here.... \(\huge\bf{y-y_{1}=m(x-x_{1})}\) We input the point in as \(\large\bf{(x_{1},y_{1})}\). \(\huge\bf{y+9=m(x-7)}\) If we simplify the slope we get \(\large\bf{\frac{2}{9}}\). We input this as m. \(\huge\bf{y+9=\frac{2}{9}(x-7)}\) This is how it would look.

OpenStudy (tanabugg):

oh it says it's wrong though

563blackghost (563blackghost):

Wait so it was 1 not 11?

OpenStudy (tanabugg):

yes

563blackghost (563blackghost):

Ok so thats why its wrong. So we have to find the slope again. \(\huge\bf{\frac{y_{2}-y_{1}}{x_{2}-x_{1}} \rightarrow \frac{1-7}{9+9}=-\frac{1}{3}}\) With this we would input to find the y-intercept by inputting point. \(\huge\bf{7=-\frac{1}{3}(-9)+b}\) Simplify. \(\huge\bf{7=3+b}\) \(\huge\bf{4=b}\) \(\huge\bf\color{red}{y=-\frac{1}{3}x+b}\)

563blackghost (563blackghost):

Sorry forgot to input... \(\huge\bf\color{red}{y=-\frac{1}{3}x+4}\)

OpenStudy (tanabugg):

what about slope intercept on the equation on theline??

563blackghost (563blackghost):

Sorry that was for the slope-intercept. The point slope is this... \(\huge\bf\color{red}{y-7=-\frac{1}{3}(x+9)}\) So so sorry.

563blackghost (563blackghost):

The point is `(7,-9)` we input into the equation as `(x1,y2)`

OpenStudy (tanabugg):

oh lol that's why I got it arong again lol

563blackghost (563blackghost):

Im so sorry

OpenStudy (tanabugg):

It'll be fine but what was the answer for slope-inter?

563blackghost (563blackghost):

\(\huge\bf\color{red}{y=-\frac{1}{3}x+4}\)

563blackghost (563blackghost):

I hope they are correct :) I must take my leave, good luck with your work ^.^

OpenStudy (tanabugg):

thanks!